OpenStudy (anonymous):

How would I set this up? Karen uses her credit card to purchase a new television for $695.20. She can pay off up to $325 per month. The card has an annual rate of 17.9% compounded monthly. How much will she pay in interest?

OpenStudy (amistre64):

Bnew = Bold k - P when Bnew = 0, its paid off, add up all the payments made and subtract off the original amount to see how much you paid for the convience of being broke ...

OpenStudy (amistre64):

I would set it up as: \[0=Bk^n-P\frac{1-k^n}{1-k}\] \[Bk^n=P\frac{1-k^n}{1-k}\] \[\frac BPk^n(1-k)=1-k^n\] \[\frac BP(1-k)=k^{-n}-1\] \[\frac BP(1-k)+1=k^{-n}\] \[log\left(\frac BP(1-k)+1\right)=-n~logk\] \[\frac{log\left(\frac BP(1-k)+1\right)}{-logk}=n\] to evaluate the number of payments

OpenStudy (amistre64):

in this case, k = 1+.179/12

OpenStudy (anonymous):

What does be stand for?

OpenStudy (amistre64):

the original balance to payoff

OpenStudy (amistre64):

it will take 2 months and a little more to pay off

OpenStudy (anonymous):

I'm not look for the months though. I need to find out how much she will pay in interest

OpenStudy (anonymous):

@manguitar I don't see how that is helpful in anyway...

OpenStudy (amistre64):

you need to know how many payments are made, and since the payments are made each month ... you need to know how many months it takes to payoff

OpenStudy (anonymous):

it is trust me 2 dollars full of chicken!

OpenStudy (anonymous):

Oh! Okay, that makes sense @amistre64

OpenStudy (amistre64):

we can work this longhand since its so short: 695.20 = starting balance first payment: 695.20(1+.179/12) - 325 = 380.57 second payment: 380.57(1+.179/12) - 325 = 61.25 second payment: 61.25(1+.179/12) = 62.16 so a total payment amount of 325 + 325 + 62.16 gives us the amount of the original balance plus interest

OpenStudy (amistre64):

lol, the last payment is a third

OpenStudy (anonymous):

and full of chicken

OpenStudy (amistre64):

712.16 is what i determine for total payments; subtract off the 695.20 to determine the excess paid in interest. the amount for "n" (2.19) that i originally determined would have given us a total payment of: 325(2.19) = 711.75 since it assumes that it doesnt take a full 3 months and ignores the remaining days of the third month. but its pretty close :)

OpenStudy (amistre64):

...and full of chickens i spose

OpenStudy (anonymous):

yes full of chickens

OpenStudy (anonymous):

Sorry for the delayed reply, I'm a little confused on that last paragraph you sent

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i was just comparing the longhanded version with the first approach i used.

OpenStudy (amistre64):

if we work out the accumulation of interest and the payments made, we get a total payout of: 325 + 325 + 62.16 = 712.16 this total payment amount covers the original balance, plus interest 712.16 = 695.20 + int paid 712.16 - 695.20 = int paid

OpenStudy (anonymous):

So, the final answer would be $16.96?

OpenStudy (amistre64):

with any luck, yes :)

OpenStudy (anonymous):

$16.96 full of chicken @AdrianK

OpenStudy (anonymous):

Thank you so much! You were really helpful! @amistre64 :D

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

don't forget full of chicken